Flora11:

Sean spends 3/5 of his money on a present for his mother. He spent 3/4 of the remainder on a present for his sister. a) What fraction of his money did he spend on the present for his sister? b) If he spent $450 altogether, how much money did he have left?

justjm:

Let us denote the money Sean originally had as m. He spends (3/4)m on his mother. He spent 3/4 of the remainder on his sister. R=m-(3/4)m=(1/4)m. He spends (3/4)R = (3/4)(1/4)m = (3/16) on his sister. If he spends $450 altogether that would mean he spent (3/5)(450) on his mother leaving $80, and 3/4 of that on his sister leaving $20.